# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ VI-D Simulation Experiment
A simulation experiment is conducted to verify the proposed state-centric necessary condition. As shown in Fig. 2(b), our previous work [8] provided a 4th order suboptimal trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, $u=u\left(t\right)$, $t\in\left[t,t_{\mathrm{f}}\right]$. The augmented switching law of the original trajectory is ${\mathcal{L}}=\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}\overline{01}\underline{0}\overline{0}$. According to Corollary 3, $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=6>4$ implies that ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is not optimal. Let ${\boldsymbol{H}}$ be the equality constraints induced by $\mathcal{L}$. Note that ${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}=9$ and $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}\in{\mathbb{R}}^{9\times 11}$ except the 4th and 11th columns is full row rank. Then, disturb $t_{4}$ into $t_{4}^{\prime}$ and search for the minimum terminal time $t_{\mathrm{f}}^{\prime}$ subject to the constraints of (53a). As shown in Fig. 2(a), $t_{\mathrm{f}}^{\prime}$ achieves local minimum $t_{\mathrm{f}}-0.0061$ at $t_{4}^{\prime}=t_{4}+0.0208$, resulting in the optimized trajectory ${\boldsymbol{x}}^{\prime}$ shown in Fig. 2(b). It can be observed in Fig. 2(b4) that $x_{3}^{\prime}$ is tangent to $-x_{{\mathrm{m}}3}$, while $x_{3}>-x_{{\mathrm{m}}3}$ holds in the original trajectory. In other words, the original trajectory fails to fully utilize the feasible set, and the extreme of ${t_{\mathrm{f}}}^{\prime}$ activates a new constraint; hence, ${\boldsymbol{t}}^{\prime}$ cannot be further optimized under fixing ${\mathcal{L}}$. Among them,
$\begin{dcases}{\boldsymbol{t}}=\left(1.7500,2.1562,3.1562,4.5342,5.5342,\right.\\ \quad\left.6.0342,6.5221,7.0100,7.2226,8.9165,9.8604\right),\\ {\boldsymbol{t}}^{\prime}=\left(1.7500,2.1562,3.1562,4.5550,5.5550,\right.\\ \quad\left.6.2023,6.6621,7.1218,7.2166,8.9104,9.8543\right).\end{dcases}$ (68)
Therefore, ${\boldsymbol{x}}^{\prime}$ achieves a $0.6\%$ reduction in the terminal time compared to ${\boldsymbol{x}}$. The proposed state-centric necessary condition is verified through the simulation experiment. More efficient optimization algorithms based on the proposed state-centric necessary condition can be developed in future works.
VII Conclusions
@@
This paper has set out to establish an innovative theoretical framework for the augmented switching law for time-optimal control of controllable linear systems and its applications. The proposed augmented switching law represents the input control and the feasibility of the Bang-Singular-Bang trajectory in a compact form. Specifically, the equality and inequality constraints induced by the augmented switching law represent a sufficient and necessary condition for the feasibility of the Bang-Singular-Bang trajectory. Based on the augmented switching law, a given feasible trajectory can be disturbed with the fixing augmented switching law and feasibility guarantees, resulting in a state-centric necessary condition for optimal control since any disturbed feasible trajectory should have a strictly longer terminal time than the optimal terminal time due to the uniqueness of optimal control. The proposed necessary condition states that the Jacobian matrix induced by the augmented switching law must not be full row rank. High order conditions can be derived based on the local optimality of the optimal trajectory. Different from the traditional necessary condition based on adjoining methods, the developed necessary condition requires only state and control information without dependence on costate information.
The proposed state-centric necessary condition is applied to the optimal control for chain-of-integrators systems with full box constraints. Under the condition that every arc is strictly feasible except for the constraints active during the whole arc, the proposed state-centric necessary condition implies that the number of arcs should be no greater than the sum of the order of constrained arcs plus the order of the problem. Considering the chattering phenomenon induced by the 2nd order state constraints, this paper proves a recursive equation for the junction time before the chattering limit point. The above two conclusions are challenging to prove by traditional costate-based necessary conditions, which demonstrates the effectiveness of the proposed state-centric necessary conditions.
@@ Acknowledgment
This work was supported by the National Key Research and Development Program of China under Grant 2023YFB4302003.

### v2

@@ VII-B Determining the Optimality of a 4th-Order Trajectory
A near-optimal trajectory planning method is proposed in our previous work [20]. Consider a 4th-order trajectory for high-order chain-of-integrator systems (48a) provided in [20] where ${\boldsymbol{x}}_{0}=\left(0.75,-0.375,2,9\right)$, ${{\boldsymbol{x}}_{\mathrm{f}}}=\left(0.25,0.5,-2,-5\right)$, ${\boldsymbol{x}}_{\mathrm{m}}=\left(1,1.5,4,20\right)$, and ${u_{\mathrm{m}}}=1$, as shown in Fig. 5(b). The ASL of the original trajectory ${\boldsymbol{x}}\left(t\right)$ is ${\mathcal{L}}=\underline{01}\overline{0}\underline{2}\overline{01}\underline{0}\overline{01}\underline{0}\overline{0}$. Note that $N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|=6>4$; hence, Corollary 2 implies that the terminal time ${t_{\mathrm{f}}}=9.8604$ is not optimal. Let ${\boldsymbol{H}}$ be the equality constraints induced by $\mathcal{L}$. Note that ${\mathrm{rank}}\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}=9$ and $\frac{\partial{\boldsymbol{H}}}{\partial{\boldsymbol{t}}}\in{\mathbb{R}}^{9\times 11}$ except the 4th and 11th columns has full row rank. Then, perturb $t_{4}$ to $t_{4}^{\prime}$ and search for the minimum terminal time $t_{\mathrm{f}}^{\prime}$ subject to the constraints of (48a). As shown in Fig. 5(a), $t_{\mathrm{f}}^{\prime}$ achieves its local minimum $t_{\mathrm{f}}-0.0061$ at $t_{4}^{\prime}=t_{4}+0.0208$, resulting in the optimized trajectory ${\boldsymbol{x}}^{\prime}$. It can be observed in Fig. 5(b2) that $x_{3}^{\prime}$ is tangent to $-x_{{\mathrm{m}}3}$, while $x_{3}>-x_{{\mathrm{m}}3}$ holds in the original trajectory. Hence, the original trajectory fails to fully utilize the feasible set, and the extreme of ${t_{\mathrm{f}}}^{\prime}$ activates a new constraint. Specifically, ${\boldsymbol{x}}^{\prime}$ achieves a 0.06% reduction in the terminal time compared to ${\boldsymbol{x}}$. Therefore, the proposed state-centric necessary condition can sensitively detect the non-optimality of a given trajectory based on state information.
VII-C Optimizing a 5th-Order Feasible BBS Trajectory
Arc-based [36] and discretized optimization [37] methods are two main approaches to solving high-order problems. Consider a 5th-order problem for high-order chain-of-integrator systems (48a) where ${\boldsymbol{x}}_{0}={\boldsymbol{0}}$, ${{\boldsymbol{x}}_{\mathrm{f}}}={\boldsymbol{e}}_{5}$, $u_{\mathrm{m}}=1$, and ${\boldsymbol{x}}_{\mathrm{m}}=\left(0.8,0.5,0.5,0.5,1\right)$. A feasible BBS trajectory ${\boldsymbol{x}}_{1}$ is planned by [36], as shown in Fig. 6(a). Denote $\mathrm{DOF}\triangleq N-\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|-n$; hence, Corollary 2 implies that $\mathrm{DOF}\leq 0$ holds for an optimal trajectory. As shown in Fig. 6(b), $\mathrm{DOF}_{1}=6>0$ which does not satisfy the proposed state-centric necessary condition. Fix the first 5 arcs as well as the ASL, and vary the motion time of the 6th arc to search the minimal terminal time until reaching the boundary of the feasible set. If some arcs reach zero time, then delete the system behavior from the ASL. Repeat the above process, resulting in trajectories ${\boldsymbol{x}}_{i}$, $i\in\left[6\right]$. As shown in Figs. 6(a) and 7(b), the terminal time decreases through iteration, and the DOF finally reaches 0 after 5 iterative steps. Finally, the optimized trajectory ${\boldsymbol{x}}_{6}$ achieves a 17.7% reduction in the terminal time compared to the original trajectory ${\boldsymbol{x}}_{1}$. In fact, it remains challenging to plan near-optimal high-order trajectories using arc-based methods despite advantages in computational accuracy and analytical simplicity. The proposed state-centric condition has the potential to optimize BBS trajectories planned by arc-based methods, which will be investigated in future works.
Figure 6: Iteration of a 5th-order feasible BBS trajectory. The initial trajectory ${\boldsymbol{x}}_{1}$ is planned by [36]. (a) The states and controls. (b) The ASLs.
A discretized optimization method [37] is applied to plan the trajectory ${\boldsymbol{x}}_{\text{base}}$ for comparison, where the time domain is discretized into 1000 intervals. An optimal solution is obtained with a 0.016% relative error of terminal time compared to ${\boldsymbol{x}}_{6}$, as shown in Fig. 7(b). Therefore, [37] can plan optimal or near-optimal trajectories for the high-order problem. However, $u_{\text{base}}$ exhibits oscillation, as shown in Fig. 7(a1), while $u_{i}$ is BBS. Defining the total variation of the control as $T_{\mathrm{v}}\triangleq\int_{0}^{{t_{\mathrm{f}}}}\left|u\left(t\right)\right|\mathrm{d}t$ to describe the trajectory stability, ${\boldsymbol{x}}_{5}$ achieves an 8.9% $T_{\mathrm{v}}$ of ${\boldsymbol{x}}_{\text{base}}$. For discretized optimization like [37], the open-loop errors of terminal states are difficult to constrain due to numerical error in each interval. Define $E_{\mathrm{s}}=\sqrt{\sum_{k=1}^{n}\left(\frac{\hat{x}_{{\mathrm{f}}k}-x_{{\mathrm{f}}k}}{x_{\mathrm{m}k}}\right)^{2}}$ where $\hat{\boldsymbol{x}}_{\mathrm{f}}=\left(\hat{x}_{{\mathrm{f}}k}\right)_{k=1}^{n}$ is directly integrated from the control input. As shown in Fig. 7(d), arc-based trajectories ${\boldsymbol{x}}_{i}$, $i\in\left[6\right]$, reduce at least 5 orders of magnitude of $E_{\mathrm{s}}$ compared to ${\boldsymbol{x}}_{\text{base}}$. Hence, an approach combining arc-based methods and the proposed state-centric necessary condition has the potential to plan near-optimal trajectories with higher trajectory stability and computational accuracy compared to discretized optimization.
Figure 7: Comparison of the original trajectory ${\boldsymbol{x}}_{1}$ [36], the optimized trajectory ${\boldsymbol{x}}_{6}$ through the proposed state-centric necessary condition, and the trajectory ${\boldsymbol{x}}_{\text{base}}$ planned by [37]. (a) The states and controls. (b) The terminal time. (c) The total variation of the control. (d) The open-loop error of terminal states.
VIII Conclusions
@@
This paper has set out to establish an innovative theoretical framework for time-optimal control of controllable linear systems with a single input. The proposed augmented switching law (ASL) represents the input control and the feasibility of a bang-bang and singular (BBS) trajectory in a compact form. Specifically, the equality and inequality constraints induced by the ASL represent a sufficient and necessary condition for the trajectory’ feasibility. Based on the ASL, a given feasible trajectory can be perturbed with feasibility guarantees, resulting in a state-centric necessary condition for time optimality since any perturbed feasible trajectory should have a strictly longer terminal time than the optimal one. The proposed necessary condition states that the Jacobian matrix induced by the ASL must not be full row rank. In contrast to traditional costate-based conditions, the developed necessary condition requires only states without dependence on costate information.
The proposed state-centric necessary condition is applied to the optimal control for chain-of-integrator systems with full box constraints from both theoretical and numerical perspectives. An upper bound of the number of arcs is determined in a general sense, and a recursive equation for the junction time in chattering induced by 2nd-order constraints is derived. Note that the above two conclusions are challenging to prove by costate-based conditions. Numerical experiments showed that the proposed state-centric necessary condition can sensitively detect the non-optimality of a given trajectory based on state information, and it has the potential to optimize feasible BBS trajectories with higher trajectory stability and computational accuracy compared to discretized optimization methods.
@@ Acknowledgment
This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFB4302003, and in part by the National Natural Science Foundation of China under Grant 624B2077.
Appendix A An Example of Application of the Proposed State-Centric Necessary Condition
Let’s introduce Corollary 4 of order $n$ as follows. We will show why it is challenging to derive the existence/non-existence of chattering based on PMP-based necessary conditions, and how to obtain the result based on the proposed state-centric necessary condition. Specifically, we will try to derive the result based on PMP-based/state-centric necessary conditions as examples.
A-A Problem Statement
Consider the following simplified 4th-order chain-of-integrator system:
$\min$ ${t_{\mathrm{f}}}$ (62)
s.t. $\dot{x}_{1}=u,\,\dot{x}_{k}=x_{k-1},\,k\in\left\{2,3,\dots,n\right\},$
$\left|u\right|\leq 1,x_{2}\leq 1,$
${\boldsymbol{x}}\left(0\right)={\boldsymbol{x}}_{0},\,{\boldsymbol{x}}\left({t_{\mathrm{f}}}\right)={\boldsymbol{x}}_{\mathrm{f}}.$
Now, the question is: Is there a $\left({\boldsymbol{x}}_{0},{\boldsymbol{x}}_{\mathrm{f}}\right)$, s.t. the optimal control is chattering? In other words, can the optimal control switch infinitely many times in a finite interval $\left[0,{t_{\mathrm{f}}}\right]$?
(For the case where $n=4$, the answer is NO according to Corollary 4 in the manuscript.)
A-B Preliminaries
To compare fairly the proposed state-centric necessary condition and PMP-based necessary conditions, we provide some preliminaries for both approaches.
Assume that chattering occurs in the optimal control. Then, $\exists\left\{t_{i}\right\}_{i\in{\mathbb{N}}}\subset\left[0,{t_{\mathrm{f}}}\right]$ increasing strictly monotonically, s.t.
  • $\left\{t_{i}\right\}_{i\in{\mathbb{N}}}$ has a finite limit point, i.e.,
    $t_{\infty}\triangleq\lim_{i\to\infty}t_{i}\in\left[0,{t_{\mathrm{f}}}\right].$ (63)
    Furthermore, the chattering (unconstrained) arcs converge to a constrained arc at $t_{\infty}$, i.e.,
    $\lim_{t\to t_{\infty}}x_{1}=0,\,\lim_{t\to t_{\infty}}x_{2}=1.$ (64)
  • $\forall i\in{\mathbb{N}}$, $x_{2}\left(t_{i}\right)=1$ and $x_{1}\left(t_{i}\right)=0$ hold.
  • $\forall i\in{\mathbb{N}}$, the control in $\left(t_{i},t_{i+1}\right)$ satisfies
    $u\left(t\right)=\begin{dcases}-1,&t\in\left(t_{i},\frac{3}{4}t_{i}+\frac{1}{4}t_{i+1}\right),\\ 1,&t\in\left(\frac{3}{4}t_{i}+\frac{1}{4}t_{i+1},\frac{1}{4}t_{i}+\frac{3}{4}t_{i+1}\right),\\ -1,&t\in\left(\frac{1}{4}t_{i}+\frac{3}{4}t_{i+1},t_{i+1}\right).\end{dcases}$ (65)
The proof of the above conclusions can be found in our related work [32] (Section IV-B).
The above conclusions are visualized in Figure 8.
Figure 8: Chattering phenomena in time-optimal control for 4th-order chain-of-integrator systems with constraints on $u$ and $x_{2}$.
A-C Why Deriving the Existence/Non-Existence of Chattering through PMP-Based Necessary Conditions is Challenging?
Now, we try to derive the existence/non-existence of chattering based on PMP-based necessary conditions, which will be shown challenging.
The Hamiltonian of problem (62) is
${\mathcal{H}}\left({\boldsymbol{x}}\left(t\right),u\left(t\right),\lambda_{0},{\boldsymbol{\lambda}}\left(t\right),{\boldsymbol{\eta}}\left(t\right),t\right)=\lambda_{0}+\lambda_{1}u+\sum_{k=2}^{n}\lambda_{k}x_{k-1}+\eta\left(x_{2}-1\right).$ (66)
The costate vector ${\boldsymbol{\lambda}}$ satisfies the Hamilton’s equations, i.e.,
$\begin{dcases}\dot{\lambda}_{2}=-\lambda_{3}+\eta,\\ \dot{\lambda}_{n}=0,\\ \dot{\lambda}_{k}=-\lambda_{k+1},\,k\in\left\{1,3,4,5,\dots,n-1\right\}.\\ \end{dcases}$ (67)
Furthermore, $\forall t\in\left[0,{t_{\mathrm{f}}}\right]$, it holds that
$\begin{dcases}\lambda_{0}\geq 0,\,\left(\lambda_{0},{\boldsymbol{\lambda}}\left(t\right)\right)\not={\boldsymbol{0}},\\ \eta\left(t\right)\geq 0,\,\eta\left(t\right)\left(x_{2}\left(t\right)-1\right)=0,\\ {\mathcal{H}}\left(t\right)\equiv 0.\end{dcases}$ (68)
PMP implies that it holds a.e. that
$u\left(t\right)=\begin{dcases}-1,&\lambda_{1}\left(t\right)>0,\\ 0,&\lambda_{1}\left(t\right)\equiv 0\text{ holds for a period},\\ 1,&\lambda_{1}\left(t\right)<0.\\ \end{dcases}$ (69)
Specifically, in a period, $u\equiv 0$ if and only if a constrained arc occurs where $x_{2}\equiv 1$.
Furthermore, a junction of the costate vector ${\boldsymbol{\lambda}}$ occurs when the state constraint $x_{2}\leq 1$ switches between active and inactive. Specifically, at junction time $t_{i}$, $i\in{\mathbb{N}}$, it holds that
$\begin{dcases}\lambda_{2}\left(t_{i}^{+}\right)\leq\lambda_{2}\left(t_{i}^{-}\right),\\ \lambda_{k}\left(t_{i}^{+}\right)=\lambda_{k}\left(t_{i}^{-}\right).\end{dcases}$ (70)
Consider the trajectory in $\left[t_{0},t_{\infty}\right]$. Denote
$\begin{dcases}\lambda_{0k}\triangleq\lambda_{k}\left(t_{0}^{+}\right),\,\forall k\in\left\{1,2,\dots,n\right\},\\ \mu_{i}=\lambda_{2}\left(t_{i}^{-}\right)-\lambda_{2}\left(t_{i}^{+}\right)\geq 0,\,\forall i\in{\mathbb{N}}^{*}.\end{dcases}$ (71)
Then costates in $\left(t_{0},t_{\infty}\right)$ are
$\begin{dcases}\lambda_{k}\left(t\right)=\sum_{j=k}^{n}\frac{\lambda_{0j}}{\left(j-k\right)!}\left(t_{0}-t\right)^{j-k},\,k\in\left\{3,4,\dots,n\right\},\\ \lambda_{2}\left(t\right)=\sum_{j=2}^{n}\frac{\lambda_{0j}}{\left(j-2\right)!}\left(t_{0}-t\right)^{j-2}-\sum_{i=1}^{\infty}\mu_{i}\delta_{\left(t_{i},t_{\infty}\right)}\left(t\right),\\ \lambda_{1}\left(t\right)=\sum_{j=1}^{n}\frac{\lambda_{0j}}{\left(j-1\right)!}\left(t_{0}-t\right)^{j-1}-\sum_{i=1}^{\infty}\mu_{i}\left(t_{i}-t\right)\delta_{\left(t_{i},t_{\infty}\right)}\left(t\right),\end{dcases}$ (72)
where
$\delta_{\left(t_{i},t_{\infty}\right)}\left(t\right)=\begin{dcases}1,&t\in\left(t_{i},t_{\infty}\right),\\ 0,&\text{otherwise}.\end{dcases}$ (73)
In other words, $\forall k\geq 3$, $\lambda_{k}\left(t\right)$ is a polynomial of order $\left(n-k-1\right)$. $\forall k\leq 2$, $\lambda_{k}\left(t\right)$ is a piecewise polynomial of order $\left(n-k-1\right)$ due to the junction condition (70).
For convenience, let
$\lambda_{1i}\left(t\right)=\sum_{j=1}^{n}\frac{\lambda_{0j}}{\left(j-1\right)!}\left(t_{0}-t\right)^{j-1}-\sum_{p=1}^{i-1}\mu_{p}\left(t_{p}-t\right),\,i\in{\mathbb{N}}^{*}.$ (74)
Then,
$\lambda_{1}\left(t\right)=\lambda_{1i}\left(t\right),\,\forall i\in{\mathbb{N}}^{*},\,t\in\left(t_{i-1},t_{i}\right).$ (75)
According to (65) and (69), it holds that
$\lambda_{1i}\left(\frac{1}{4}t_{i}+\frac{3}{4}t_{i-1}\right)=\lambda_{1i}\left(\frac{3}{4}t_{i}+\frac{1}{4}t_{i-1}\right)=0,\,\forall i\in{\mathbb{N}}.$ (76)
In other words, $\forall i\in{\mathbb{N}}$, $\exists\hat{t}_{i}\not\in\left[t_{i-1},t_{i}\right]$, s.t.
$\lambda_{1i}\left(t\right)=\left(t-\frac{1}{4}t_{i}-\frac{3}{4}t_{i-1}\right)\left(t-\frac{3}{4}t_{i}-\frac{1}{4}t_{i-1}\right)p_{i}\left(t\right),$ (77)
where $p_{i}$ is a polynomial of order no higher than $n-3$. Furthermore,
$\forall t\in\left(t_{i-1},t_{i}\right),\,p_{i}\left(t\right)>0.$ (78)
Therefore, a sufficient condition for the non-existence of chattering is $\mathcal{C}=\varnothing$, where
$\mathcal{C}=\left\{\begin{aligned} &\left(\left(\mu_{i}\right)_{i\in{\mathbb{N}}^{*}},\left(t_{i}\right)_{i\in{\mathbb{N}}},\left(p_{i}\left(\bullet\right)\right)_{i\in{\mathbb{N}}},\left(\lambda_{1i}\left(\bullet\right)\right)_{i\in{\mathbb{N}}}\right):\\ &\qquad t_{i}<t_{i+1},\,\lim_{i\to\infty}t_{i}<\infty,\,\lambda_{0n}\not=0,\,\mu_{i}\geq 0,\\ &\qquad p_{i}\left(\bullet\right)\text{ is a polynomial of order no higher than }n-3,\\ &\qquad\forall t\in\left(t_{i-1},t_{i}\right),\,p_{i}\left(t\right)>0,\\ &\qquad\forall t\in{\mathbb{R}},\,\lambda_{1i}\left(t\right)=\lambda_{0n}\frac{\left(-1\right)^{n-1}}{\left(n-1\right)!}\left(t-\frac{1}{4}t_{i}-\frac{3}{4}t_{i-1}\right)\left(t-\frac{3}{4}t_{i}-\frac{1}{4}t_{i-1}\right)p_{i}\left(t\right),\\ &\qquad\forall t\in{\mathbb{R}},\,\lambda_{1\,i+1}\left(t\right)=\lambda_{1i}\left(t\right)+\mu_{i}\left(t-t_{i}\right)\end{aligned}\right\}.$ (79)
It is significantly challenging to prove that $\mathcal{C}_{\pm}=\varnothing$. A key step is deriving a recursive expression for $\left(t_{i}\right)_{i\in{\mathbb{N}}}$. However, the above key step requires eliminating the polynomial $p_{i}$, which involves more complex computations. As far as we have tried, it is too complex to obtain the recursive expression for $\left(t_{i}\right)_{i\in{\mathbb{N}}}$ based on (79). In contrast, our proposed Corollary 4 can directly derive the recursive expression for $\left(t_{i}\right)_{i\in{\mathbb{N}}}$ in a compact and simple form.
A-D Why Deriving the Existence/Non-Existence of Chattering through the Proposed State-Centric Necessary Condition is Possible?
Denote
$\tau_{i}=t_{\infty}-t_{i}.$ (80)
Then, chattering is allowed only if $\tau_{0}\in{\mathbb{R}}_{++}$ and $\lim_{i\to\infty}\tau_{i}=0$.
According to Corollary 4, the recursive expression for $\left(\tau_{i}\right)_{i\in{\mathbb{N}}^{*}}$ is
$\det\left(\left(\tau_{i-k}+3\tau_{i-k-1}\right)^{j+1}-3\left(3\tau_{i-k}+\tau_{i-k-1}\right)^{j+1}+3\left(\tau_{i-k+1}+3\tau_{i-k}\right)^{j+1}-\left(3\tau_{i-k+1}+\tau_{i-k}\right)^{j+1}\right)_{j\in\left[n-2\right],k\in\left[n-2\right]}=0.$ (81)
Once the initial value $\left(\tau_{i}\right)_{i=1}^{n-2}$ is given, the whole sequence $\left(\tau_{i}\right)_{i\in\infty}$ is determined. In other words, a sufficient condition for the non-existence of chattering is $\widehat{\mathcal{C}}=\varnothing$, where
[…]
Evidently, the condition on (82) is much simpler than the condition on (79). Specifically, the condition on (82) requires analyzing the convergence of a series for the recursive expressions (81) in the form of a system of polynomial equations.
A-E Deriving the Non-Existence of Chattering through the Proposed State-Centric Necessary Condition in 4th-Order Problem
Based on the proposed state-centric necessary condition, we can even directly prove that $\widehat{\mathcal{C}}=\varnothing$ when $n=4$; hence, chattering does not exist in the optimal control of problem (62). Note that the above claim is a novel theoretical result.
Let $n=4$. Then, the recursive expression (81) for $\left(\tau_{i}\right)_{i\in{\mathbb{N}}^{*}}$ is
$\det\left(\left(\tau_{i-k}+3\tau_{i-k-1}\right)^{j+1}-3\left(3\tau_{i-k}+\tau_{i-k-1}\right)^{j+1}+3\left(\tau_{i-k+1}+3\tau_{i-k}\right)^{j+1}-\left(3\tau_{i-k+1}+\tau_{i-k}\right)^{j+1}\right)_{j\in\left[2\right],k\in\left[2\right]}$ (83)
$=$ $144\left(\tau_{i-1}-\tau_{i-3}\right)\left(\tau_{i}-\tau_{i-2}\right)f\left(\tau_{i-1}-\tau_{i};\tau_{i-2}-\tau_{i-1},\tau_{i-3}-\tau_{i-2}\right)=0,$
where
$f\left(z;z_{1},z_{2}\right)=\left(z_{1}-z_{2}\right)z^{2}+\left(z_{1}^{2}+z_{1}z_{2}-z_{2}^{2}\right)z-z_{2}\left(z_{1}^{2}+z_{1}z_{2}-z_{2}^{2}\right).$ (84)
Note that $\tau_{i}<\tau_{i-1}<\tau_{i-2}<\tau_{i-3}$. Let $z_{i}=\tau_{i-1}-\tau_{i}>0$. Then,
$f\left(z_{i};z_{i-1},z_{i-2}\right)=0.$ (85)
$\lim_{i\to\infty}\tau_{i}=0$ implies that $\lim_{i\to\infty}z_{i}=0$. Without loss of generality, let $z_{1}>z_{2}$. Then,
$\begin{dcases}f\left(z_{2};z_{1},z_{2}\right)=\left(z_{1}-z_{2}\right)^{2}z_{2}>0,\\ f\left(0;z_{1},z_{2}\right)=-z_{2}\left(z_{1}^{2}+z_{1}z_{2}-z_{2}^{2}\right)<0.\end{dcases}$ (86)
Note that $f\left(z;z_{1},z_{2}\right)$ is a quadratic function of $z$. Therefore, $f\left(z;z_{1},z_{2}\right)=0$ has two roots in $\left(-\infty,0\right)$ and $\left(0,z_{2}\right)$, respectively. Since $z_{3}>0$ and $f\left(z_{3};z_{1},z_{2}\right)=0$, it holds that $0<z_{3}<z_{2}<z_{1}$. For the same reason recursively, it holds that
$0<z_{i}<z_{i-1},\,\forall i\geq 2$ (87)
Let $r_{i}=1-\frac{z_{i+1}}{z_{i}}\in\left(0,1\right)$. Then, (85) implies that
$g\left(r_{i+1},r_{i}\right)\triangleq r_{i+1}^{2}-\left(3+\frac{1}{r_{i}\left(1-r_{i}\right)}\right)r_{i+1}+1=0.$ (88)
Note that
$\begin{dcases}g\left(0,r_{i}\right)=1>0,\\ g\left(r_{i},r_{i}\right)=-\frac{r_{i}\left(2-r_{i}\right)^{2}}{1-r_{i}}<0,\\ g\left(1,r_{i}\right)=-1-\frac{1}{r_{i}\left(1-r_{i}\right)}<0.\end{dcases}$ (89)
(88) has two roots in $\left(0,r_{i}\right)$ and $\left(1,+\infty\right)$, respectively. Hence, $10<r_{i+1}<r_{i}<$ 1. Denote $r_{\infty}\triangleq\lim_{i\to\infty}r_{i}\in\left[0,1\right]$. Then, $g\left(r_{\infty},r_{\infty}\right)$ implies that $r_{\infty}=0$. Note that
$r_{i+1}=\frac{1}{2}\left[-\left(3+\frac{1}{r_{i}\left(1-r_{i}\right)}\right)+\sqrt{\left(3+\frac{1}{r_{i}\left(1-r_{i}\right)}\right)^{2}-4}\right]=r_{i}-4r_{i}^{2}+\mathcal{O}\left(r_{i}^{3}\right),\,i\to\infty.$ (90)
By Stolz-Cesàro theorem, it holds that
$\lim_{i\to\infty}i\left(\frac{z_{i}}{z_{i+1}}-1\right)=\lim_{i\to\infty}\frac{ir_{i}}{1-r_{i}}=\lim_{i\to\infty}\frac{i}{\frac{1}{r_{i}}}=\lim_{i\to\infty}\frac{i+1-i}{\frac{1}{r_{i+1}}-\frac{1}{r_{i}}}=\lim_{i\to\infty}\frac{r_{i}^{2}\left(1+\mathcal{O}\left(r_{i}\right)\right)}{4r_{i}^{2}+\mathcal{O}\left(r_{i}^{3}\right)}=\frac{1}{4}\in\left(0,1\right).$ (91)
By Raabe-Duhamel’s test, it holds that
$\infty=\sum_{i=1}^{\infty}z_{i}=\tau_{0}-\tau_{\infty}=t_{\infty}-t_{0}<\infty,$ (92)
which is a contradiction.
Therefore, the set $\widehat{\mathcal{C}}$ defined in (82) is empty. Therefore, chattering does not exist in the optimal control of problem (62) when $n=4$.
A-F Summary
In this appendix, we provide an example to compare the proposed state-centric necessary condition and PMP-based necessary conditions.
  1. 1.
    We show that it is challenging to derive the existence/non-existence of chattering based on PMP-based necessary conditions in the considered optimal control problem. The key step is to derive the recursive expression for junction times, which involves complex computations.
  2. 2.
    We show that the proposed state-centric necessary condition can directly derive the recursive expression for junction times in a compact and simple form, which is much simpler than the condition based on PMP-based necessary conditions. Specifically, we prove that chattering does not exist in the considered optimal control problem of order 4.